Calculate the magnitude of the electric field at one corner of a square 2.12 m on a side if the other three corners are occupied by 6.75×10−6 C charges.
Answer #1

Electric field due to the charge
= kq/r^2 = (9x10^9) * (6.75 x 10^-6) / (2.12)^2 N/C = 13516.83


The resultant of two such charges are
= 2cos45° * [ 13516.83] N/C = 19115.69 N/C

Electric field due to charge
= kq/(√2r)^2 = (9x10^9) * (6.75 x 10^-6) / (2.12√2)^2 = 6758.42 N

Total field = 19115.69+ 6758.42 = 25874.100 N/C